OpenStudy (anonymous):

two trains start from towns 144 mi apart and travel towards eachother on parallel tracks they pass eachother 1.6 hour later if one train travels 10 mph faster then the other find the speed of each train

OpenStudy (dumbcow):

distance = rate*time the combined distance is 144 for both trains and the time it takes is 1.6 hrs \[1.6r_1 + 1.6r_2 = 144\] and \[r_1 = r_2 +10\] substitute and find r_1 and r_2

OpenStudy (dumbcow):

you could also rearrange the equations -by dividing by 1.6 \[r_1 + r_2 = 90\] and \[r_1 -r_2 = 10\] so what 2 numbers add to 90 but have a difference of 10?

OpenStudy (anonymous):

40 and 50

OpenStudy (dumbcow):

yep...those are the speeds of 2 trains

OpenStudy (anonymous):

thank you so much!

OpenStudy (dumbcow):

yw...i hope you understood the equations because most of the time the solution isn't this straightforward and you need to solve the system with substitution or elimination

OpenStudy (anonymous):

yes i understand im horrible at math and i have been having a hard time picking this up

OpenStudy (dumbcow):

keep practicing, once you understand the best way to get good at math is practice and repitition